OpenStudy (anonymous):

Given the function g(x) = -4x + 5, compare and contrast g(2) and g(-4). Choose the statement that is true concerning these two values. A The values of g(2) and g(-4) cannot be compared. B The value of g(2) is the same as the value of g(-4). C The value of g(2) is larger than the value of g(-4). D The value of g(2) is smaller than the value of g(-4).

OpenStudy (hoblos):

g(2) = -3 (-4) = 21 The value of g(2) is smaller than the value of g(-4)

OpenStudy (anonymous):

Well. g(2) = -4(2) + 5= -3 g(-4)= -4(-4) +5 = 21 21> -3 g(-4) > g(2)